# TCP-Diffusion: A Multi-modal Diffusion Model for Global Tropical Cyclone Precipitation Forecasting with Change Awareness

### v1

@@ Introduction
where $\theta$ represents all the trainable parameters in our model. $r_{s}$ could be calculated by Equation 4.
For the inference phase, we consider a random noise $r$ as the $D_{y}^{N}$ (Ho, Jain, and Abbeel 2020) and use the following Equation 7 to denoise $D_{y}^{N}$ step by step to obtain the $\hat{D}_{y}^{0}$, which is equivalent to $\hat{D}_{y}$:
@@
where $\hat{r}_{s}$ is the predicted noise added at $s$-th step, $\sigma_{s}$ is a variance hyperparameter. $\epsilon\sim\mathcal{N}(0,1)$ is a random noise, which plays a critical role in generating high-quality prediction in the inference phase (Ho, Jain, and Abbeel 2020).
Environment Awareness 3DUNet (EA-3DUNet)
EA-3DUNet is the main component of TCP-Diffusion. It can extract various features from multi-modal data sources. Due to the differing characteristics and dimensions of these heterogeneous meteorological data, we need to build multiple encoders to embed these data. Therefore, we build the Historical Data Encoder for Rainfall Data and Environment Data, and the Future Prediction Data Encoder for ECMWF’s Future prediction data encoding. To better extract temporal and spatial information from these data, we build 3DUNet with a temporal and spatial attention mechanism. We will introduce all these modules in the following sections.
Historical Data Encoder.
Rainfall Data $X_{rain}$, adjacent residual data $D_{x}$, ERA5 surface data $X_{SfEnv}$, and ERA5 pressure level data $X_{PlEnv}$ are two-dimensional (2D) data. Considering that our data has a time dimension, we build Historical $Data_{2d}$ Encoder with 3D Convolutional Neural Networks (CNNs) to encode all 2D data with time information first. To reduce computational resources, we use one module to encode these data, so we concatenate them to get $X_{his2D}=[X_{rain},D_{x},X_{SfEnv},X_{PLEnv},r_{s}]$. This process is defined as follows:
@@
where $W_{his2D}$ is the parameter of $Conv3d$ (Historical $Data_{2d}$ Encoder). $e_{his2D}$ represents the features encoded from Historical 2D Data $X_{his2D}$ and serves as the initial feature input to 3DUNet for further feature extraction.
The scalar TC variable $X_{InEnv}$ does not have a 2D structure, so we build Historical $Data_{1d}$ Encoder with Multilayer Perceptron (MLP) and Transformer (Vaswani et al. 2017) layers to encode the $X_{InEnv}$ and obtain the time information. The main process is as follows:
$e_{mlp}=\phi(X_{InEnv},W_{mlp})$ (9)
$e_{his1D}=Transformer(e_{mlp},W_{atten})$ (10)
where $\phi$ is denoted as an MLP module, which is used for getting the features $e_{mlp}$ of different variables. $W_{mlp}$ is the parameter of $\phi$. $Transformer$ module is used for obtaining the time information of $X_{InEnv}$. $W_{atten}$ is the parameter of $Transformer$ and $e_{his1D}$ is the encoding feature of $X_{InEnv}$ at the initial stage of our model.
Future Prediction Data Encoder.
Future Prediction Data $X_{future}$ contain future information and some physical rules provided by NWP, which differ from Historical data. It is challenging for a single module to encode these data containing different types of information at the same time (Alzubaidi et al. 2021). Therefore, we build a Future $Data_{2d}$ Encoder to encode these data and try to obtain more specific features. We use a modified Resnet-18 (He et al. 2016) to obtain the vector. The main process of Future $Data_{2d}$ Encoder is as follows:
$e_{future}=Resnet(X_{future},W_{res18})$ (11)
where $Resnet$ is the Resnet-18 model we used, $W_{res18}$ is denoted as the parameters of Resnet-18, and $e_{future}$ is the condition extracted from $X_{future}$ and used for guide our model to make a better prediction.
@@
3DUNet (Çiçek et al. 2016) is the core component of EA-3DUNet and is a classical DL structure for tasks involving 2D data with time information. It usually includes three parts shown in Figure 2: U-Net encoder ($UNet_{en}$), U-Net decoder ($UNet_{de}$), and the bottleneck between encoder and decoder. There are several modules in $UNet_{en}$ and $UNet_{de}$, called $module_{en}$ and $module_{dn}$ respectively. In each module, we build different blocks (blue cuboids) to capture features from various aspects.
Specifically, in the encoder module ($module_{en}$), shown as the blue cuboids in $UNet_{en}$ in Figure 2, we stack two CNN blocks, one spatial attention (SA) block, one temporal attention (TA) block, and a down-sampling block. The network progressively performs downsampling by stacking several $module_{en}$. Each $module_{en}$ receives the feature map from the previous module and also the conditions $e_{his1D}$ and $e_{future}$ from the Historical $Data_{1d}$ Encoder and Future $Data_{2d}$ Encoder, respectively. Additionally, step $s$ is also received to inform the model how much noise to remove at the current stage. We concatenate these conditions to form the final condition $Cond=[e_{his1D},e_{future},s]$. The main process of $UNet_{en}$ is as follows:
where $module_{en}^{i}$ is the depth-$i$ module in $UNet_{en}$, $i\in\{1,2,\ldots,K\}$. $K$ is a hyper-parameter. $W_{en_{i}}$ is the parameters of the $module_{en}^{i}$. if $module_{en}^{i}$ is the first one $(i=1)$, $module_{en}^{1}$ receives the $e_{his2D}$ from historical data encoder. if $module_{en}^{i}$ is not the first one $(i>1)$, $module_{en}^{i}$ receives the $e_{en}^{i-1}$ from $module_{en}^{i-1}$.
The module between $UNet_{en}$ and $UNet_{de}$ is the Bottleneck. The Bottleneck also contains two CNN blocks, one SA block, and one TA block. The main process is as follows:
$e_{mid}=Bottleneck(e_{en}^{K},Cond,W_{neck})$ (13)
For $UNet_{de}$, its structure is similar to that of $UNet_{en}$. They both include $K$ modules with similar architecture. In $module_{de}$, there are two CNN blocks, one SA block, one TA block and the final up-sampling block. There are also skip connections between $UNet_{en}$ and $UNet_{de}$. Thus, the definition of $UNet_{de}$ is as follows:
where $module_{de}^{i}$ is the depth-$i$ module in $UNet_{de}$, $i\in\{K,\ldots,2,1\}$. $W_{de_{i}}$ is the parameters of the $module_{de}^{i}$. if $module_{de}^{i}$ is the deepest one $(i=K)$, $module_{de}^{K}$ receives the $e_{mid}$ from the Bottleneck. if $module_{de}^{i}$ is not the deepest one $(i<K)$, $module_{de}^{i}$ receives the $e_{de}^{i+1}$ from $module_{de}^{i+1}$. if $i=1$, $e_{de}^{i}$ is the final output of EA-3DUNet, which means $\hat{r}_{s}=e_{de}^{1}$. Then, we could calculate the loss between $r_{s}$ and $\hat{r}_{s}$ to optimize the parameters of our model.
Figure 3: The prediction results of different DL methods on the TC DUMAZILE in North Indian Ocean at 03/03/2018 06:00. The first column is the previous 4 timesteps of rainfall data that is used as input. The second column is the future TC rainfall we want to predict. Each subsequent column from the left shows the predictions from a different DL method. “TCP-Diff #1” and “TCP-Diff #2” are the results from TCP-Diffusion with different initial random noise $r$.
Experiments
Evaluation metrics
@@
The Equitable Threat Score (ETS) (Gandin and Murphy 1992) is a metric used to evaluate the accuracy of precipitation forecasts. Compared to the more popular metric, the Critical Success Index (CSI) (Schaefer 1990), it considers the effects of random chance, providing a more equitable assessment of a forecast model’s performance (Manzato and Jolliffe 2017). The definition of ETS can be checked in Supplementary Material.
We use this metric to show the performance for predicting light, medium, and heavy rainfall, setting thresholds 6 mm/3hr (ETS-6), 24 mm/3hr (ETS-24), and 60 mm/3hr (ETS-60) respectively.
@@
TPMAE is a critical index for representing the intensity of TC rainfall. We calculate the absolute difference between real TC rainfall $Y$ and the predicted rainfall $\hat{Y}$, which is denoted as Total Precipitation Mean Absolute Error (TPMAE, mm/3hr), to show the prediction skill of different methods at predicting total TC rainfall.
Comparison with State-of-the-art DL Methods
TCP-Diffusion is compared with a deterministic DL method: U-Net (Çiçek et al. 2016). We also compare our model with two generative DL methods developed for nowcasting precipitation: PreDiff, which is DM-based (Gao et al. 2024), and NowcastNet (Zhang et al. 2023). We also compare TCP-Diffusion with a basic baseline, persistence forecasting, using the last observed rainfall field for each future prediction (Panofsky 1963).
@@
Figure 3 shows an example of a single prediction from TCP-Diffusion and comparison DL methods. Comparing the classical rainfall prediction model, like U-Net, the DM-based models (Predif and TCP-Diffusion) can give more realistic rainfall predictions with fine spatial detail, such as the shape of the rain band. U-Net and NowcastNet produce predictions that are too spatially smooth. This is one of the reasons we chose DMs as the basis for our TC rainfall prediction model. When comparing our TCP-Diffusion with PreDiff in this example, our model provides more accurate results and does better at predicting the increasing trend of precipitation intensity with time, because the ARP mechanism gives our model the capability to track the rainfall change. We show two samples from TCP-Diffusion, generated using different noise inputs. Overall, they have similar large-scale rainfall trends and structures. This means that the information extracted from other input data, such as historical TC rainfall and 2D environmental variables, can guide our model to make reasonable and realistic forecasts. Overall, TCP-Diffusion can produce rainfall predictions that have realistic spatial and temporal structure, performing better than the other DL methods.
@@
Table 1: Comparison with other SOTA DL methods. Equitable Threat Score (ETS) is used to evaluate the prediction skill of each model. ETS-6, ETS-24, and ETS-60 are used to show the prediction skill for light ($>$6 mm/3hr), medium ($>$24 mm/3hr), and high ($>$60 mm/3hr) rainfall respectively. TPMAE shows the total precipitation forecasting skill of different methods and, with unit mm/3hr.
To further demonstrate the effectiveness of our method, we calculate and compare various metrics of our method with other methods on the entire test set, as shown in Table 1. To ensure the stability and reliability of the results, we randomly select half of the samples from the entire test set, calculate the metrics 10 times, and obtain the mean of these 10 results. For light rainfall prediction (ETS-6), U-Net achieves the best performance, with our method ranking second. However, the performance of U-Net on the other three metrics is unsatisfactory because light rainfall constitutes the majority of TC rainfall. This causes U-Net to focus on the accuracy of light rainfall and neglect the accuracy of moderate and heavy rainfall predictions, which is why it performs well in light rainfall prediction but poorly in heavy rainfall prediction. For higher rainfall thresholds and TPMAE, our method achieves the best performance. Especially in heavy rainfall prediction, TCP-Diffusion shows significant improvement compared to other DL models. It is also the only model to perform better than Persistence forecasting, indicating the value of designing a specialised system for the task of TC rainfall prediction.
@@
Figure 4: TC rainfall forecasting distributions of different DL methods. The grey histogram is the distribution of MSWEP observations. The coloured lines show histograms of rainfall rates from different DL forecasting methods. The circular magnified region more clearly shows differences between these methods in the high-frequency span of rainfall intensity. Note the logarithmic vertical axis.
Figure 4 shows histograms of rainfall rates for observations and forecasts by each DL method. We find that the performance of DM-based models, PreDiff and TCP-Diffusion, is much better than that of non-DM-based models, especially in producing a realistic frequency of the heaviest rainfall rates. Our method predicts a slightly higher frequency of the heaviest precipitation intensities than is observed, while PreDiff predicts a slightly lower precipitation intensity. However, for more moderate rainfall intensities, TCP-Diffusion has a more realistic frequency than PreDiff, as shown in the circled region in Figure 4.
Comparison with NWP Methods

### v2

@@ 3.2 Diffusion Model
where $\theta$ represents all the trainable parameters in our model. $r_{s}$ is calculated by Equation 4.
For the inference phase, we consider a random noise $r$ as the $D_{y}^{N}$ (Ho et al., 2020) and use the following Equation 7 to denoise $D_{y}^{N}$ step by step to obtain the $\hat{D}_{y}^{0}$, which is equivalent to $\hat{D}_{y}$:
@@
where $\hat{r}_{s}$ is the predicted noise added at $s$-th step, $\sigma_{s}$ is a variance hyperparameter. $\epsilon\sim\mathcal{N}(0,1)$ is a random noise, which plays a critical role in generating high-quality prediction in the inference phase (Ho et al., 2020).
Overall, the training phase (Figure 2.(b)) and inference phase (Figure 2.(a) and (c)) can be represented by the following pseudocode. Here, $q(\cdot)$ denotes the data distribution from which samples are drawn. Specifically, $q(D_{y})$ represents the distribution of the target variable, while $q(X_{history})$ and $q(X_{future})$ represent the distributions of historical and future input features, respectively.
Algorithm 1 Training
1:  repeat
2:     $D_{y}^{0}\sim q(D_{y})$
3:     $s\sim\text{Uniform}(\{1,\dots,N\})$
4:     $D_{y}^{s}=\sqrt{\bar{\alpha_{s}}}D_{y}^{0}+\sqrt{1-\bar{\alpha_{s}}}r$
5:     $X_{history}\sim q(X_{history})$
6:     $X_{future}\sim q(X_{future})$
7:     $r\sim\mathcal{N}(0,1)$
8:     $\hat{r}=EA3DUnet(X_{history},X_{future},D_{y}^{s},s,r)$
9:     Take gradient descent step on $\nabla_{\theta}\lVert r-\hat{r}\rVert^{2}$
10:  until converged
Algorithm 2 Inference
1:  $D_{y}^{N}\sim\mathcal{N}(0,1)$
2:  $X_{history}\sim q(X_{history})$
3:  $X_{future}\sim q(X_{future})$
4:  for $s=N,\dots,1$ do
5:     $\epsilon\sim\mathcal{N}(0,1)$ if $s>1$, else $\epsilon=0$
6:     $\hat{r}=EA3DUnet(X_{history},X_{future},\hat{D}_{y}^{s},s,r)$
7:     $\hat{D}_{y}^{s-1}=\frac{1}{\sqrt{\alpha_{s}}}(\hat{D}_{y}^{s}-\frac{\beta_{s}}{\sqrt{1-\bar{\alpha_{s}}}}\hat{r})+\sigma_{s}\epsilon$
8:  end for
9:  Use Equation.3 to get $\hat{Y}$
10:  return $\hat{Y}$
3.3 Environmentally-Aware 3DUNet (EA-3DUNet)
EA-3DUNet is the neural network component of TCP-Diffusion. It can extract various features from multi-modal data sources. Due to the differing characteristics and dimensions of these heterogeneous meteorological data, we need to build multiple encoders to embed these data. Therefore, we build the Historical Data Encoder for Rainfall Data and Environment Data, and the Future Prediction Data Encoder for future prediction data. To better extract temporal and spatial information from these data, we build 3DUNet with a temporal and spatial attention mechanism. We will introduce all these modules in the following sections.
Historical Data Encoder
Rainfall Data $X_{rain}$, adjacent residual data $D_{x}$, ERA5 surface data $X_{SfEnv}$, and ERA5 pressure level data $X_{PlEnv}$ are two-dimensional (2D) data. Considering that our data has a time dimension, we build a Historical $Data_{2d}$ Encoder with 3D Convolutional Neural Networks (CNNs) to encode all 2D data with time information first. To reduce computational resources, we use one module to encode these data, so we concatenate them to get $X_{his2D}=[X_{rain},D_{x},X_{SfEnv},X_{PLEnv},r_{s}]$. This process is defined as follows:
@@ 3.3 Environmentally-Aware 3DUNet (EA-3DUNet)
where $W_{his2D}$ is the encoder parameters. $e_{his2D}$ represents the features encoded from $X_{his2D}$ and serves as the initial feature input to 3DUNet for further feature extraction.
The scalar TC variable $X_{Sc}$ does not have a 2D structure, so we build a Historical $Data_{1d}$ Encoder with Multilayer Perceptron (MLP) and Transformer (Vaswani et al., 2017) layers to encode $X_{Sc}$ and obtain the temporal information. The main process is as follows:
$e_{mlp}=\phi(X_{Sc},W_{mlp})$ (9)
$e_{his1D}=\mathrm{Transformer}(e_{mlp},W_{atten})$ (10)
where $\phi$ is denoted as an MLP module, which is used for getting the features $e_{mlp}$ of different variables. $W_{mlp}$ is the parameters of $\phi$. The $\mathrm{Transformer}$ module is used for obtaining the temporal information of $X_{Sc}$. $W_{atten}$ is the parameter of $\mathrm{Transformer}$ and $e_{his1D}$ is the encoded features of $X_{Sc}$.
Future Prediction Data Encoder
Future prediction data $X_{future}$ contain future information and some physical information provided by NWP, which differs from Historical data. It is challenging for a single module to encode these data containing different types of information at the same time (Alzubaidi et al., 2021). Therefore, we build a Future $Data_{2d}$ Encoder to encode these data and try to obtain more specific features. We use a modified Resnet-18 (He et al., 2016) to obtain the vector. The main process of Future $Data_{2d}$ Encoder is as follows:
$e_{future}=\mathrm{Resnet}(X_{future},W_{res18})$ (11)
where $\mathrm{Resnet}$ is the Resnet-18 model, $W_{res18}$ is the parameters, and $e_{future}$ is the condition extracted from $X_{future}$ and used for guiding our model to make a better prediction.
@@
3DUNet (Çiçek et al., 2016) is the core component of EA-3DUNet and is a classical DL structure for tasks involving 2D data with time information. It usually includes three parts shown in Figure 2.(b): U-Net encoder ($U\text{-}Net_{en}$), U-Net decoder ($U\text{-}Net_{de}$), and the bottleneck between encoder and decoder. There are several modules in $U\text{-}Net_{en}$ and $U\text{-}Net_{de}$, called $module_{en}$ and $module_{de}$ respectively. In each module, we build different blocks (blue cuboids) to capture features from various aspects. Over all, the 3DUNet receives the feature map from the Historical $Data_{1d}$ Encoder and the condition information from Future $Data_{2d}$ Encoder. Additionally, step $s$ is also included in the condition information to inform the model how much noise to remove at the current stage. Then, the 3DUNet predicts the noise ($\hat{r}_{s}$) added at the step $s$. Finally, there are two pathes shown in Figure 2.(b). In the training stage (green arrow), we could calculate the loss between actual noise $r_{s}$ and $\hat{r}_{s}$ to optimize the parameters of our model. In the inference stage (red arrow), $\hat{r}_{s}$ could be used for obtaining $D^{s-1}_{y}$ using Equation 7. The main process of 3DUNet is shown in Appendix.B.
Figure 3: The prediction results of different DL methods on TC Dumazile in the North Indian Ocean at 03/03/2018 06:00. The first column is the previous 4 timesteps of rainfall data used as input. The second column is the future TC rainfall we want to predict. Each subsequent column from the left shows the predictions from a different DL method. “TCP-Diff #1” and “TCP-Diff #2” are two samples from our TCP-Diffusion model with different initial random noise $r$.
4 Dataset and Evaluation Methods
4.1 Dataset
We collected various TC rainfall-related data for our TCP-Diffusion model to comprehensively represent TC rainfall. A total of 1877 TCs spanning from 1980 to 2020 are collected, covering the six major ocean areas. These TC data are divided into three sets: training set (1751 TCs), validation set (87 TCs), and test set (126 TCs from 2018 to 2020). We divided the data into two parts: $X_{historical}$ and $X_{future}$. Details of the data we used are in Appendix.A.
4.2 Evaluation Metrics
@@
The Equitable Threat Score (ETS) (Gandin & Murphy, 1992) is a metric used to evaluate the accuracy of precipitation forecasts. Compared to the Critical Success Index (CSI) (Schaefer, 1990), a more popular metric, it considers the effects of random chance, providing a more equitable assessment of a forecast model’s performance (Manzato & Jolliffe, 2017). The definition of ETS is given in Appendix.C. We use this metric to show the performance for predicting light, medium, and heavy rainfall, setting thresholds 6 mm/3hr (ETS-6), 24 mm/3hr (ETS-24), and 60 mm/3hr (ETS-60) respectively.
@@
: This is the absolute difference between real TC rainfall $Y$ and the predicted rainfall $\hat{Y}$ averaged over the 10 spatial domain, to show the prediction skill of different methods at predicting total TC rainfall.
5 Results
5.1 Comparison with State-of-the-art DL Methods
TCP-Diffusion is compared with a deterministic DL method: U-Net (Çiçek et al., 2016). We also compare our model with two generative DL methods developed for nowcasting precipitation: PreDiff, which is DM-based (Gao et al., 2024), and NowcastNet (Zhang et al., 2023). We also compare TCP-Diffusion with a basic baseline, persistence forecasting, using the last observed rainfall field for each future prediction (Panofsky, 1963). Since TCP-Diffusion is a probabilistic prediction model, to ensure the stability of the results presented, we conducted eight tests on the test set. For each test, the evaluation metrics were calculated, and the final results were obtained by averaging the metrics across these eight tests. Additionally, the individual results from each test are provided in Table 5 in Appendix.D.2.
@@ 5.1 Comparison with State-of-the-art DL Methods
Figure 3 shows an example of predictions from TCP-Diffusion and other DL methods for a forecast of TC Dumazile. Comparing with the deterministic U-Net rainfall prediction model, the DM-based models (Predif and TCP-Diffusion) can give more realistic rainfall predictions with fine spatial detail, such as the shape of the rain band. U-Net and NowcastNet produce predictions that are too spatially smooth. This is one of the reasons we chose DMs as the basis for our TC rainfall prediction model. When comparing our TCP-Diffusion with PreDiff in this example, our model provides more accurate results and does better at predicting the increasing trend of precipitation intensity with time, because the ARP mechanism gives our model the capability to track the rainfall change. We show two samples from TCP-Diffusion, generated using different noise inputs. We also show all 8 samples from TCP-Diffusion in Figure 9 in Appendix.D. Overall, they have similar large-scale rainfall trends and structures. This means that the information extracted from other input data, such as historical TC rainfall and 2D environmental variables, can guide our model to make reasonable and realistic forecasts. Overall, these results indicate that TCP-Diffusion can produce rainfall predictions that have realistic spatial and temporal structure, performing better in this example than the other DL methods.
@@
Table 1: Comparison with other SOTA DL methods. Equitable Threat Score (ETS) is used to evaluate the prediction skill of each model. ETS-6, ETS-24, and ETS-60 are used to show the prediction skill for light ($>$6 mm/3hr), medium ($>$24 mm/3hr), and high ($>$60 mm/3hr) rainfall respectively. TPMAE shows the total precipitation forecasting skill of different methods and, with unit mm/3hr. The results here represent the overall performance of each model for lead times of 3, 6, 9, and 12 hours.
To further demonstrate the effectiveness of our method, we calculate and compare various metrics on the entire test set, as shown in Table 1. We also show the performance of TCP-Diffusion compared with other methods at lead times of 3, 6, 9, and 12 hours separately in Figure 8 in Appendix.D. The lead time indicates the amount of time between the issuance of a forecast and the actual occurrence of the weather event. For light rainfall prediction (ETS-6), U-Net achieves the best performance, with TCP-Diffusion ranking second. However, U-Net performs poorly for moderate and heavy rainfall predictions. For higher rainfall thresholds and TPMAE, our TCP-Diffusion method achieves the best performance. Especially in heavy rainfall prediction, TCP-Diffusion shows a substantial improvement compared to other DL models. It is also the only model that performs better than Persistence forecasting, indicating the value of designing a specialised system for the task of TC rainfall prediction. We also show some analysis about why Persistence forecasting performs better than the other three DL models in Appendix.D.
@@
Figure 4: TC rainfall frequency distributions of different DL forecasting methods. The grey histogram is the distribution of MSWEP observations. The coloured lines show histograms of rainfall rates from different DL forecasting methods. The circular magnified region more clearly shows differences between these methods in the indicated span of rainfall intensity. Note the logarithmic vertical axis.
Figure 4 shows histograms of rainfall rates for observations and forecasts by each DL method. We find that the performance of DM-based models, PreDiff and TCP-Diffusion, is much better than that of non-DM-based models, especially in producing a realistic frequency of the heaviest rainfall rates. Our method predicts a slightly higher frequency of the heaviest precipitation intensities than is observed, while PreDiff predicts a slightly lower precipitation intensity. However, for more moderate rainfall intensities, TCP-Diffusion has a more realistic frequency than PreDiff, as shown in the circled region in Figure 4. We also present the distribution of rainfall prediction from TCP-Diffusion and compared methods at each lead-time in Figure 10 in Appendix.D.
Analysis of TC Rainfall’s Radially Averaged Power Spectral Density
We use the Radially Averaged Power Spectral Density (RAPSD) to help demonstrate that our model better captures the spatial structure of large-scale TC rainfall systems, as shown in Figure 5. If the power spectral density curve predicted by the model is closer to the ground truth MSWEP at a given spatial frequency, it indicates that the model captures the rainfall variability at that spatial scale more effectively. The low wavenumber region corresponds to large-scale weather systems, such as fronts and tropical cyclones. As observed in Figure 5, the power spectral density of TCP-Diffusion is closer to MSWEP compared to PreDiff in the low wavenumber region, indicating that our method achieves better capture of rainfall characteristics at the TC scale. We also show the RAPSD of rainfall prediction from TCP-Diffusion and compared methods at each lead-time in Figure 11 in Appendix.D.
Figure 5: Radially Averaged Power Spectral Density (RAPSD) analysis across datasets. The plots compare the power spectral density of predicted rainfall using different models over a range of wavenumber. In the low wavenumber region, higher power spectral density represents that the displayed data belongs to large-scale weather systems, such as fronts and tropical cyclones. The power spectral density of TCP-Diffusion in the low wavenumber region is closer to the ground truth rainfall (MSWEP), indicating that our method achieves better capture of rainfall characteristics at the TC scale. Note the logarithmic vertical axis.
5.2 Comparison with NWP Methods